Can u help with variation?
Describe the combined variation that is modeled by each formula.
V = Bh÷3
V = (r^2)h
h = V÷(r^2)
------------
Write the funtion that models each variation. Find z when x = 4 and y = 9.
z varies directly with x and inversely with y. When x = 6 and y = 2, z = 15.
z varies jointly with x and y. When x = 2 and y = 3, z = 60
z varies directly with the square of x and inversely with y. When x = 2 and y = 4, z = 3.

Answer by stanbon(48558)

Describe the combined variation that is modeled by each formula.
V = Bh÷3
V varies directly with B; V varies directly with h; B and h vary inversely
----------------
V = (r^2)h
V varies directly with r^2; V varies directly with h; r^2 and h vary inversely
------------
h = V÷(r^2)
h varies directly with V; h varies inversely with r^2; V and r^2 vary directly
------------
Write the funtion that models each variation.
Find z when x = 4 and y = 9.
z varies directly with x and inversely with y.
z = kx/y ; z = (4/9)k (k is the constant of variation)
---------------
When x = 6 and y = 2, z = 15.
z varies jointly with x and y.
z = kxy
15 = (6*2)k
15 = 12k
k=5/4
-------------

When x = 2 and y = 3, z = 60
z varies directly with the square of x and inversely with y.
z = kx^2/y
60 = k(4/3)
k=45
--------------
Cheers,
Stan H.


When x = 2 and y = 4, z = 3.